OpenStudy (anonymous):

What is the value of b if –7 + b = –8?

OpenStudy (anonymous):

-7+b=-8 +7 +7 __________________ b=-1

OpenStudy (anonymous):

Thank you every much

OpenStudy (anonymous):

ahahahha ok allright